You currently have 20,000X ethidium bromide. You want to make 250 mL of 1X ethidium bromide dissolved in 1X TBE. You currently also have a stock of 10X TBE solution. How much 10X TBE do you need and how much 20,000X ethidium bromide do you need?
Answer #1

Here we want to make a solution of 250ml containing 1x ethidium bromide and 1x TBE

Let's calculate how much ethidium we will need

We have 20000 ethidium bromide

By using the simple formula

C1V1=C2V2

20000× V1 = 1×250

V1= 250/20000

V1=0.0125 ml

So0.0125 ml of ethidium bromide will be needed

similarly for TBE

C1V1=C2V2

10 × V1= 1×250

V1= 250/10

V1= 2.5 ml

So2.5 ml of 10x TBE will be needed

so to make the above mentioned solution of 250 ml 0.0125 ml of 20000x ethidium bromide and 2.5 ml of 10x TBE will be needed.

i hope this helps if you have any query or want more detailed explanation fill Free to ask in the comment section below.